In the quark model of fundamental particles, a proton is composed of three quarks: two "up" quarks, each having charge +2e/3, and one "down" quark, having charge -e/3. Suppose that the three quarks are equidistant from one another. Take the distance to be 1.48×10-15 m and calculate the potential energy of the subsystem of two "up" quarks. (MeV)

Calculate the total electric potential energy of the three-particle system. (MeV)
Answer #1

potential energy of the subsystem of two "up" quarks = Kq1*q2/r^2

= 9*10^9*(2/3*1.6*10^-19)^2/(1.48*10^-15)^2 = 291478352614132 = 2.91*10^14 Mev

total electric potential energy of the three-particle system=

Kq1*q2/r^2 - Kq1*q3/r2^2 -kq2*q3/r3^2

= 2.91*10^14 - (2.91*10^14 )/8 - (2.91*10^14 )/2

= 1.09*10^14 Mev